# Error threshold in active steering protocols for few-qubit systems

### v1

@@ III.1 One system qubit
On the other hand, one reaches the target state with high purity for $\Gamma\ll\Gamma_{c}$. In this case, errors remain correctable in our protocol. For $\Gamma>\Gamma_{c}$, however, the dynamics is dominated by dissipation channels associated with $\Gamma$ in Eq. (II.1.1). We observe that one approaches a pure state again, see also Fig. 3(b), where the purity $P$ increases towards its maximum value $P=1$ as $\Gamma$ increases. For large $\Gamma$, this pure state is linked to the dark state of the dissipative Lindbladian instead of the target state $|\Psi_{f}\rangle$ which for high amplitude damping and dephasing is given by the ground state $|1\rangle$ of $H_{s}$. Remarkably, our numerical simulations show that for $\Gamma\approx\Gamma_{c}$, the competition between the unitary time evolution due to steering operators and the measurement- and noise-induced dissipative processes results in an infinite-temperature state, where control over the system state is basically lost and the purity gap closes.
In order to rationalize the existence of a threshold transition when varying the error rate $\Gamma$, let us consider the measurement-averaged state change $d\overline{\rho}$ in one time step, starting from the actual state $\rho=\rho(t)$. Neglecting $E_{A}\ll J$ in $H_{s}$ and writing the steering operator as $H_{s,\alpha}=\pm J\sigma^{\alpha}$, Eq. (II.1.1) yields
$d\overline{\rho}$ $=$ $\mp iJ\delta t\,[\sigma^{\alpha},\rho]+\delta t\Gamma_{\rm ms}\,{\cal D}[\sigma^{s}]\rho$ (29)
@@
Here three different energy scales appear, namely $J$, $\Gamma$ (assuming $\Gamma_{\rm AD}=\Gamma_{\rm PD})$, and $\Gamma_{\rm ms}=\Lambda^{2}\delta t$. For $J\ll\Gamma_{\rm ms}$, a balance between measurement-induced and noise-induced terms is expected for $\Gamma_{c}\approx\Gamma_{\rm ms}$. However, this balance does not result in a purity gap closing transition. In fact, in this limit, our numerical simulations do not find a sharp threshold transition as described above. For the case $J\gg\Gamma_{\rm ms}$ considered above, on the other hand, the existence of the error threshold is robust under changes of $J$ and $\Lambda$. In this regime, the steering dynamics is clearly important, and one has to consider the competition between the unitary term and the damping terms (due to $\Gamma_{\rm ms}$ and $\Gamma$) as well as the active decision making.
From Eq. (29), assuming $O_{s}=\sigma^{z}$, we can analytically compute the anticipated average change $\overline{\langle dO_{s}\rangle_{\alpha}}={\rm Tr}(d\overline{\rho}\,O_{s})$ from Eqs. (11) and (29). Writing $\rho(t)=\frac{1}{2}(\sigma^{0}+{\bf r}\cdot\bm{\sigma})$ in terms of a Bloch vector ${\bf r}={\bf r}(t)=(r_{x},r_{y},r_{z})^{T}$, we find
$\overline{\langle dO_{s}\rangle_{\alpha}}$ $=$ ±$\pm 2J\delta t\sum_{j}\varepsilon_{z\alpha j}r_{j}-\delta t\,\Gamma_{\rm AD}(1+r_{z})$
@@
$\overline{\langle dO_{s}\rangle_{\alpha}}=2\delta t\,{\bf r}\cdot\left(\begin{array}[]{c}\mp J\delta_{\alpha,y}\\ \pm J\delta_{\alpha,x}\\ -\Gamma_{\rm ms}\end{array}\right),$ (32)
@@ III.2 Two system qubits
Again assuming that both amplitude and phase noise are present, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$, the main panel of Fig. 4 shows that for weak error rate $\Gamma$, the $N=2$ protocol is again able to converge the time-evolving state $\rho(t)$ with high fidelity towards a predesignated target state $|\Psi_{f}\rangle$. In this specific example, we start from the initial state $|\Psi(0)\rangle=|++\rangle$ and steer towards the symmetric Bell target state, $|\Psi_{f}\rangle=|\Phi_{0,+}\rangle$, see Eq. (25). We observe that the protocol is able to retain high fidelity for arbitrarily long times, such that the state is autonomously stabilized. For an individual measurement trajectory, as shown in the main panel of Fig. 4, small-amplitude Rabi oscillations can be resolved in the time-dependent fidelity. The frequency of these oscillations corresponds to the energy difference $2E_{A}$ between the system qubit states, see Eq. (13). As for the $N=1$ case (cf. inset of Fig. 4), quantum jumps occasionally occur and reduce the fidelity for a very short time span. However, the fidelity recovers quickly to a high value again, demonstrating the robustness of the autonomous state stabilization protocol. In fact, for weak damping, we find that the $N=2$ protocol is more efficient than its $N=1$ variant because of the state-purifying capabilities of the non-unitary jump operators for $N=2$.
Figure 7: Averaged late-time fidelity $\overline{F_{|\Phi_{0,+}\rangle}}$ vs $\Gamma/\Delta$ for the $N=2$ active steering protocol, with the initial state $|\Psi(0)\rangle=|++\rangle$ and the Bell target state $|\Psi_{f}\rangle=|\Phi_{0,+}\rangle$. We use the same parameters as in Fig. 3 for both system-plus-detector subsystems with $\Lambda/\Delta=0.49$ and different $\Delta\delta t$. We include both amplitude and phase noise, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. Note the logarithmic scale for $\Gamma/\Delta$. The threshold value $\Gamma_{c}$ is marked by an arrow. Results have been obtained by averaging over 500 measurement trajectories.
@@
For the same $N=2$ protocol, we show the averaged fidelity reached at late times as function of $\Gamma/\Delta$ in Fig. 7. Similar to the $N=1$ case in Fig. 5, we again find a sharply defined error threshold, $\Gamma=\Gamma_{c}$, separating a weak-damping regime ($\Gamma<\Gamma_{c}$) and a strong-damping regime where steering fails ($\Gamma>\Gamma_{c}$). The threshold is again accompanied by a closure of the purity gap, as shown in the inset of Fig. 6. For the considered parameters, the threshold error rate is $\Gamma_{c}/\Delta\simeq 0.15$. For $\Gamma=\Gamma_{c}$, we have an infinite-temperature state, $\overline{\rho}=\frac{1}{4}\sigma_{1}^{0}\sigma_{2}^{0}$, with minimal purity $P=1/4$. For $\Gamma\ll\Gamma_{c}$, the averaged state approaches the pure target state $|\Psi_{f}\rangle$ with high fidelity. By analogy to the QEC case, one may expect that in the limit of large $N$, the fidelity and the purity approach their ideal values for all $\Gamma<\Gamma_{c}$. However, for $N=2$, the fidelity and the purity are gradually degraded upon increasing $\Gamma$ towards $\Gamma_{c}$. For strong damping $\Gamma\gg\Gamma_{c}$, on the other hand, one approaches a pure dark state associated with the dissipator $\propto\Gamma$ in Eq. (II.2). As seen in Fig. 7, the scaling property (i.e., the averaged fidelity depends on $\Gamma/\Delta$ without explicit dependence on $\delta t$) holds again.
@@ IV Conclusions and Outlook
In this work, we have formulated active steering protocols in the presence of error channels for few-qubit systems. Our numerical simulations for $N=1$ and $N=2$ system qubits reveal a purity gap closing threshold at a specific (non-universal) error rate $\Gamma_{c}$ in such an autonomous state stabilization protocol. For $\Gamma\ll\Gamma_{c}$, the protocol steers the system to the predesignated target state with high fidelity and high purity. For $\Gamma>\Gamma_{c}$, it is not possible to reach the target state anymore, and for $\Gamma\gg\Gamma_{c}$, the (undesired) stationary state of the pure error dynamics is instead attained. The competition between the two dynamical resources – active steering and errors – each separately stabilizing a different stationary state, leads to a transition which is found to be sharp, and accompanied by a closure of the purity gap. We expect that such thresholds are likely to appear also in larger-$N$ variants of our protocol.
The approach presented here is based on an active steering protocol where error channels are accounted for within the SME. A cost function maximizing the expectation value of a physical observable has been used to guide the measurement trajectory. Since the present formulation requires that one keeps track of the system state, it is not scalable and restricted to relatively small values of $N$, see also Ref. [40].
@@
A further intriguing direction aims to generalize the target state from a single state (i.e. a single point on the Bloch sphere) to a two-dimensional state manifold defining a general qubit in order to exhibit control over the whole qubit space. This will provide an active steering variant of quantum error correction protocols [25, 27, 28]. Such protocols are not restricted to stabilizer codes.

### v2

@@ III.1 One system qubit
On the other hand, one reaches the target state with high purity for $\Gamma\ll\Gamma_{c}$. In this case, errors remain correctable in our protocol. For $\Gamma>\Gamma_{c}$, however, the dynamics is dominated by dissipation channels associated with $\Gamma$ in Eq. (II.1.1). We observe that one approaches a pure state again, see also Fig. 3(b), where the purity $P$ increases towards its maximum value $P=1$ as $\Gamma$ increases. For large $\Gamma$, this pure state is linked to the dark state of the dissipative Lindbladian instead of the target state $|\Psi_{f}\rangle$ which for high amplitude damping and dephasing is given by the ground state $|1\rangle$ of $H_{s}$. Remarkably, our numerical simulations show that for $\Gamma\approx\Gamma_{c}$, the competition between the unitary time evolution due to steering operators and the measurement- and noise-induced dissipative processes results in an infinite-temperature state, where control over the system state is basically lost and the purity gap closes. We note that the error threshold has been discussed above for the ensemble averaged system, using both the fidelity and the purity as diagnostic measures. For individual measurement trajectories, signatures of the error threshold can also be observed but these are then subject to random fluctuations, see Fig. 5(c).
In order to rationalize the existence of a threshold transition when varying the error rate $\Gamma$, let us consider the measurement-averaged state change $\langle d\rho\rangle_{\rm ms}$ in one time step, starting from the actual state $\rho=\rho(t)$. Neglecting $E_{A}\ll J$ in $H_{s}$ and writing the steering operator as $H_{s,\alpha}=\pm J\sigma^{\alpha}$, Eq. (II.1.1) yields
$\langle d\rho\rangle_{\rm ms}$ $=$ $\mp iJ\delta t\,[\sigma^{\alpha},\rho]+\delta t\Gamma_{\rm ms}\,{\cal D}[\sigma^{s}]\rho$ (29)
@@
Here three different energy scales appear, namely $J$, $\Gamma$ (assuming $\Gamma_{\rm AD}=\Gamma_{\rm PD})$, and $\Gamma_{\rm ms}=\Lambda^{2}\delta t$. The latter scale is the measurement-induced decay rate. For $J\ll\Gamma_{\rm ms}$, a balance between measurement-induced and noise-induced terms is expected for $\Gamma_{c}\approx\Gamma_{\rm ms}$. However, this balance does not result in a purity gap closing transition. In fact, in this limit, our numerical simulations do not find a sharp threshold transition as described above. For the case $J\gg\Gamma_{\rm ms}$ considered above, on the other hand, the existence of the error threshold is robust under changes of $J$ and $\Lambda$. In this regime, the steering dynamics is clearly important, and one has to consider the competition between the unitary term and the damping terms (due to $\Gamma_{\rm ms}$ and $\Gamma$) as well as the active decision making.
From Eq. (29), assuming $O_{s}=\sigma^{z}$, we can analytically compute the anticipated average change $\langle(dO_{s})_{\alpha}\rangle_{\rm ms}={\rm Tr}(\langle d\rho\rangle_{\rm ms}\,O_{s})$ from Eqs. (11) and (29). Writing $\rho(t)=\frac{1}{2}(\sigma^{0}+{\bf r}\cdot\bm{\sigma})$ in terms of a Bloch vector ${\bf r}={\bf r}(t)=(r_{x},r_{y},r_{z})^{T}$, we find
$\langle(dO_{s})_{\alpha}\rangle_{\rm ms}$ $=$ ±$\pm 2J\delta t\sum_{j}\varepsilon_{z\alpha j}r_{j}-\delta t\,\Gamma_{\rm AD}(1+r_{z})$
@@
$\langle(dO_{s})_{\alpha}\rangle_{\rm ms}=2\delta t\,{\bf r}\cdot\left(\begin{array}[]{c}\mp J\delta_{\alpha,y}\\ \pm J\delta_{\alpha,x}\\ -\Gamma_{\rm ms}\end{array}\right),$ (32)
@@
Figure 7: Active steering results for $N=2$, with the same parameters as in Fig. 3 for both system-plus-detector subsystems but using $\Lambda/\Delta=0.49$ and different $\Delta\delta t$. We include both amplitude and phase noise, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$. Note the logarithmic scales for $\Gamma/\Delta$. Results are averaged over 500 measurement trajectories. Panel (a): Averaged late-time fidelity $\overline{F_{|\Phi_{0,+}\rangle}}$ vs $\Gamma/\Delta$ for $|\Psi(0)\rangle=|++\rangle$ and the Bell target state $|\Psi_{f}\rangle=|\Phi_{0,+}\rangle$. The threshold value $\Gamma_{c}$ is marked by an arrow. Panel (b): Purity $P$ vs $\Gamma/\Delta$ for $\Delta\delta t=0.03$.
@@ III.2 Two system qubits
Again assuming that both amplitude and phase noise are present, $\Gamma_{\rm AD}=\Gamma_{\rm PD}=\Gamma$, Fig. 4(b) shows that for weak error rate $\Gamma$, the $N=2$ protocol is able to converge the time-evolving state $\rho(t)$ with high fidelity towards a predesignated target state $|\Psi_{f}\rangle$. In this specific example, we start from the initial state $|\Psi(0)\rangle=|++\rangle$ and steer towards the symmetric Bell target state, $|\Psi_{f}\rangle=|\Phi_{0,+}\rangle$, see Eq. (25). We observe that the protocol is able to retain high fidelity for arbitrarily long times, such that the state is autonomously stabilized. For an individual measurement trajectory, see Fig. 4(b), small-amplitude Rabi oscillations can be resolved in the time-dependent fidelity. The frequency of these oscillations corresponds to the energy difference $2E_{A}$ between the system qubit states, see Eq. (13). As for the $N=1$ case in Fig. 4(a), quantum jumps occasionally occur and reduce the fidelity for a very short time span. However, the fidelity recovers quickly to a high value again, demonstrating the robustness of the autonomous state stabilization protocol. In fact, for weak damping, we find that the $N=2$ protocol is more efficient than its $N=1$ variant because of the state-purifying capabilities of the non-unitary jump operators for $N=2$.
@@
For the same $N=2$ protocol, we show the averaged fidelity reached at late times as function of $\Gamma/\Delta$ in Fig. 7(a). Similar to the $N=1$ case in Fig. 5, we again find a sharply defined error threshold, $\Gamma=\Gamma_{c}$, separating a weak-damping regime ($\Gamma<\Gamma_{c}$) and a strong-damping regime where steering fails ($\Gamma>\Gamma_{c}$). The threshold is again accompanied by a closure of the purity gap, as shown in Fig. 7(b). For the considered parameters, the threshold error rate is $\Gamma_{c}/\Delta\simeq 0.15$. For $\Gamma=\Gamma_{c}$, we have an infinite-temperature state, $\overline{\rho}=\frac{1}{4}\sigma_{1}^{0}\sigma_{2}^{0}$, with minimal purity $P=1/4$. For $\Gamma\ll\Gamma_{c}$, the averaged state approaches the pure target state $|\Psi_{f}\rangle$ with high fidelity. By analogy to the QEC case, one may expect that in the limit of large $N$, the fidelity and the purity approach their ideal values for all $\Gamma<\Gamma_{c}$. However, for $N=2$, the fidelity and the purity are gradually degraded upon increasing $\Gamma$ towards $\Gamma_{c}$. For strong damping $\Gamma\gg\Gamma_{c}$, on the other hand, one approaches a pure dark state associated with the dissipator $\propto\Gamma$ in Eq. (II.2). As seen in Fig. 7(a), the scaling property (i.e., the averaged fidelity depends on $\Gamma/\Delta$ without explicit dependence on $\delta t$) holds again.
@@ IV Conclusions and Outlook
In this work, we have formulated active steering protocols in the presence of error channels for few-qubit systems. Our numerical simulations for $N=1$ and $N=2$ system qubits reveal a purity gap closing threshold at a specific error rate $\Gamma_{c}$ in such an autonomous state stabilization protocol. For $\Gamma\ll\Gamma_{c}$, the protocol steers the system to the predesignated target state with high fidelity and high purity. For $\Gamma>\Gamma_{c}$, it is not possible to reach the target state anymore, and for $\Gamma\gg\Gamma_{c}$, the (undesired) stationary state of the pure error dynamics is instead attained. The competition between the two dynamical resources – active steering and errors – each separately stabilizing a different stationary state, leads to a transition which is found to be sharp, and accompanied by a closure of the purity gap. We expect that such thresholds are likely to appear also in larger-$N$ variants of our protocol. We have presented results for amplitude and phase noise with equal rates, $\Gamma_{\rm AD}=\Gamma_{\rm PD}$, and for phase noise only, $\Gamma_{\rm AD}=0$. We note in passing that for $\Gamma_{\rm PD}=0$, i.e., if only amplitude noise is present, the numerically observed threshold behavior (not shown here) is observed to be qualitatively similar as for $\Gamma_{\rm AD}=\Gamma_{\rm PD}$.
It is worth noting that feedback control of a single qubit based on continuous measurements has a long history in quantum optics and quantum information science, see, e.g., Refs. [97, 19, 98]. Similarly, purification protocols for single qubit states are specific examples for such protocols for this type of feedback control, where the measurement basis changes through the course of the time evolution in order to maximize the state purity. Beyond the single-qubit case, certain multi-qubit entangled states can also be stabilized via feedback based on the measurement of collective system observables, see, e.g., Ref. [99]. In the present case, a weak measurement variant of entanglement swapping has been employed for two system qubits to achieve this goal.
The approach presented here is based on an active steering protocol where error channels are accounted for within the SME. A cost function maximizing the expectation value of a physical observable has been used to guide the measurement trajectory. Since the present formulation requires that one keeps track of the system state, it is not scalable and restricted to relatively small values of $N$, see also Ref. [40]. The state tracking requirement is a well-known and general obstacle for many state-based feedback schemes. However, for small and very well characterized systems, it may be possible to formulate an active steering protocol only in terms of measurable quantities. In such cases, feedback protocols could avoid state tracking. A detailed study of this interesting question is, however, beyond the scope of this paper.
@@
In this work, we have not studied the effects of errors that can induce leakage out of the computational qubit space, e.g., excitations to higher energy levels or quasiparticle poisoning effects. It will be interesting to clarify in future research whether such errors are also associated with a threshold. In Refs. [61, 65], the Andreev state dynamics has been studied allowing for parity transitions. Incorporating such transitions into our protocol as an additional error channel could be the object of future work. A further intriguing direction aims to generalize the target state from a single state (i.e. a single point on the Bloch sphere) to a two-dimensional state manifold defining a general qubit in order to exhibit control over the whole qubit space. This will provide an active steering variant of quantum error correction protocols [25, 27, 28]. Such protocols are not restricted to stabilizer codes.
All data underlying the figures presented in this work can be retrieved at zenodo: https://zenodo.org/records/11581356.